on monday mr. toad bought 10 gourmet flies and 4 gourmet worms for $34. on tuesday he bought 6 gourmet flies and 12 gourmet worms for $30. find the price of a gourmet fly and the price of a gourmet worm.(use a 5-step solution with two variables.) please help me with this question!!!!!!
Answer by dabanfield(803)

on monday mr. toad bought 10 gourmet flies and 4 gourmet worms for $34. on tuesday he bought 6 gourmet flies and 12 gourmet worms for $30. find the price of a gourmet fly and the price of a gourmet worm.(use a 5-step solution with two variables.) please help me with this question!!!!!!
Let f be the cost of each fly.
Let w be the cost of each worm.
Then:
10*f + 4*w =34
6*f + 12*w = 30
To eliminate the variable f multiply the first equation by 6 and the second equation by 10:
60*f + 24*w = 204
60*f + 120*w = 300
Subtract the second equation from the first:
(60-60)*f + (24-120)*w = 204-300
-96*w = -96
w = 1
Substitute w=1 in either of the first two equations to find the value for f.